# Mamba base PKD for efficient knowledge compression

### v1

@@ Abstract.
Knowledge Distillation, Mamba Architecture, Image Classification, Machine Learning, Computer Vision
copyright: none
1. Introduction
A preliminary version of this work was presented as a short poster titled “Mamba-PKD: A Framework for Efficient and Scalable Model Compression in Image Classification” at The 40th ACM/SIGAPP Symposium on Applied Computing (SAC ’25)  [^1] A preliminary version of this work appeared as a conference poster: https://doi.org/10.1145/3672608.3707887. This extended version includes additional methodological details, experiments, and analysis., This version extends the previous work by providing a detailed description of the methodology, additional experiments, and a more comprehensive discussion of results.
@@ 1. Introduction
Related to enhancing efficiency, new approaches based on State Space Models (SSMs) make significant contributions. These mathematical frameworks are ideal for modeling systems whose state evolves over time based on its previous states and current inputs, by considering simple matrix multiplications (Gu et al., 2021). SSMs serve as a foundation for capturing temporal dependencies, in tasks such as pixel sequences and time series in computer vision (Nguyen et al., 2022; Lei et al., 2024). However, traditional SSMs represent the entire system, treating all incoming information equally important, making them effective for modeling data but impractical for large sequences. Gu and Dao (2023) handled this inefficiency by introducing innovations in an architecture called Mamba. The central concept of this architecture is a Selective-State Space Model (S-SSM), which process information based on the current input, focusing on relevant information while discarding irrelevant data. This selective process reduces unnecessary computation, ensuring that resources are allocated to meaningful tasks.
Figure 1. Teacher-student framework for knowledge distillation, where knowledge can be transferred based on teacher-student outputs, features, and pure relations.
@@ 2. Related Work
2.1. Knowledge Distillation
Knowledge distillation has been extensively studied as a model compression technique that allows a smaller student model to learn from a larger teacher model (Gou et al., 2021; Han et al., 2015) (see Figure 1). Traditional KD includes transferring the knowledge to a student model from a teacher model’s output (Chen et al., 2017), intermediate layers (Romero et al., 2014), or only the relationships between different layers and data samples (Yim et al., 2017; Passban et al., 2021). These techniques lead to computational savings while preserving high accuracy. However, one limitation of traditional KD is the performance drop when there is a large capacity gap between the teacher and student models because the student cannot effectively represent key features of the teacher. This limitation has been addressed by various methods (Romero et al., 2014; Yim et al., 2017; Zhang et al., 2019; Passban et al., 2021) that modify distillation algorithms or employ multi-stage learning.
@@ 2.1. Knowledge Distillation
2.2. Progressive Knowledge Distillation
Building on traditional distillation methods, PKD incrementally transfers knowledge to improve the efficiency and scalability of the distillation process. In (Dennis et al., 2023), the authors proposed the B-DISTIL algorithm, which decomposes a large teacher model into an ensemble of smaller student models, each capable of refining predictions as more models are evaluated (see Figure 2). This way of distillation enables a flexible trade-off between inference time and accuracy, making it particularly effective for on-device inference, where computational resources are limited.
@@ 2.2. Progressive Knowledge Distillation
Figure 2. Progressive Knowledge Distillation, the knowledge from a complex architecture teacher is split among different simple architecture students, who are trained to progressively improve a coarse prediction.
2.3. State Space Models and Selectivity
@@
2.4. Dynamic Neural Networks and Hardware-aware Parallelism
@@ 3. Theory and Methodology
3.1. Knowledge Distillation Foundation
@@
(1) $L_{KD}=\alpha L_{hard}+(1-\alpha)L_{soft}$
@@
At a round $t$ (looking for weak learner), the algorithm maintains two probability matrices as $K^{+}_{t}\in\mathbb{R}^{N\times L}$ and $K^{-}_{t}\in\mathbb{R}^{N\times L}$, where $N$ is the number of data samples and $L$ is the number of labels. These matrices store the probabilities of positive and negative residual errors (differences between student and teacher predictions).
@@
(3) $\min_{\theta}\Big{(}-\frac{1}{\gamma}\sum_{i,j}I_{ij}^{+}\log\Big{(}1+\frac{l(x_{i})_{j}}{2B}\Big{)}+(1-I_{ij}^{+})\log\Big{(}1-\frac{l(x_{i})_{j}}{2B}\Big{)}\Big{)}$
@@
3.2. Mamba Architecture Theory
@@
To make this process more efficient, as shown in (Gu et al., 2021), the SSM can be expressed as a convolution. By unrolling the recursion, the output $y_{k}$ becomes a weighted sum of the current and previous inputs, where these weights are determined by powers of the matrix $\bm{\overline{A}}$ This leads to the construction of the convolution kernel $\bm{\overline{K}}$, which can be defined as:
(6) $\bm{\overline{K}}$ $=(\bm{C}\bm{\overline{B}},\bm{C}\bm{\overline{A}}\bm{\overline{B}},\dots,\bm{C}\bm{\overline{A}}^{k}\bm{\overline{B}},\dots)$
(7) $y$ $=x\ast\bm{\overline{K}}$
@@
For large sequences, the SSM processes all incoming data uniformly; this can lead to inefficiencies when much of the input data is irrelevant. Also, the representation depicted in equation 6 relies on the assumption of working in a Linear Time Invariant (LTI) System, which is not always true in complex contexts. Mamba introduces Selective-State-Space Models S6 that improve traditional SSMs by selectively processing only the most relevant parts of the input sequence. The selectivity is based on the discretization method, where the continuous $(A,B)$ are transformed to their discrete version $\overline{A_{k}},\overline{B_{k}}$ through $f_{A}(\Delta_{k},A)$ and $f_{B}(\Delta_{k},A,B)$. Rules as the zero-order hold (ZOH) defined in the following expresions can be used as $f_{A},f_{B}$
(8) $\bm{\overline{A_{k}}}=\exp(\Delta_{k}\bm{A})\qquad$
(9) $\bm{\overline{B_{k}}}=(\bm{A})^{-1}(\exp(\Delta_{k}\bm{A})-\bm{I})\cdot\bm{B}$
@@
Figure 3. Selective SSM, the flow of the data and operations, requires a hardware-aware algorithm that computes the hidden state $h$ in more efficient levels of the GPU memory hierarchy. The S-SSM is an LTVS that projects the input $x$ to construct a time dependence in $\Delta_{t},B_{t},C_{t}$. This process ensures selectivity during discretization ($A\rightarrow\overline{A},B\rightarrow\overline{B}$) and resetting the hidden state through $C_{t}$ to the final output $y_{t}$.
@@
Figure 4. Mamba Blocks inside Progressive Knowledge Distillation
3.3. Combining PKD and Mamba Architecture Methodology
@@ 4.1. Datasets
4.2. Knowledge Distillation Experiment Setup
@@ 5.1. Computational Efficiency
Table 1. MNIST overview of model configurations and some performance metrics, including the number of blocks, state dimension, total parameters, FLOPs, and accuracy for each model variant.
@@
Table 2. CIFAR-10 overview of model configurations and some performance metrics, including the number of blocks, state dimension, total parameters, FLOPs, and accuracy for each model variant.
@@
Figure 5. Accuracy vs. Number of Students
@@
Figure 6. Performance Teacher Students Comparison
5.2. Model Performance
@@
6. Conclusion and Future Work
@@
[…]
Acknowledgements.

### v2

@@ Abstract
1ITS Lab, Institute of Computer Science, University of Tartu, Estonia
2Litis, Université de Rouen, France
3Litis, INSA de Rouen, France
{joseluis,hadachi}@ut.ee, paul.honeine@univ-rouen.fr, abdelaziz.bensrhair@insa-rouen.fr
Keywords: Knowledge Distillation, Mamba Architecture, Image Classification, Machine Learning, Computer Vision.
1 Introduction
A preliminary version of this work was presented as a short poster titled “Mamba-PKD: A Framework for Efficient and Scalable Model Compression in Image Classification” at The 40th ACM/SIGAPP Symposium on Applied Computing (SAC ’25)  [^1] A preliminary version of this work appeared as a conference poster: https://doi.org/10.1145/3672608.3707887, This version extends the previous work by providing a detailed description of the methodology, additional experiments, and a more comprehensive discussion of results.
@@ 1 Introduction
Related to enhancing efficiency, new approaches based on State Space Models (SSMs) make significant contributions. These mathematical frameworks are ideal for modeling systems whose state evolves over time based on its previous states and current inputs, by considering simple matrix multiplications Gu et al. (2021). SSMs serve as a foundation for capturing temporal dependencies, in tasks such as pixel sequences and time series in computer vision (Nguyen et al., 2022; Lei et al., 2024). However, traditional SSMs represent the entire system, treating all incoming information equally important, making them effective for modeling data but impractical for large sequences. Gu and Dao (2023) handled this inefficiency by introducing innovations in an architecture called Mamba. The central concept of this architecture is a Selective-State Space Model (S-SSM), which process information based on the current input, focusing on relevant information while discarding irrelevant data. This selective process reduces unnecessary computation, ensuring that resources are allocated to meaningful tasks.
Figure 1: Teacher-student framework for knowledge distillation, where knowledge can be transferred based on teacher-student outputs, features, and pure relations.
@@ 2 Related Work
2.1 Knowledge Distillation
Knowledge distillation has been extensively studied as a model compression technique that allows a smaller student model to learn from a larger teacher model Gou et al. (2021); Han et al. (2015) (see Figure 1). Traditional KD includes transferring the knowledge to a student model from a teacher model’s output (Chen et al., 2017), intermediate layers (Romero et al., 2014), or only the relationships between different layers and data samples (Yim et al., 2017; Passban et al., 2021). These techniques lead to computational savings while preserving high accuracy. However, one limitation of traditional KD is the performance drop when there is a large capacity gap between the teacher and student models because the student cannot effectively represent key features of the teacher. This limitation has been addressed by various methods (Romero et al., 2014; Yim et al., 2017; Zhang et al., 2019; Passban et al., 2021) that modify distillation algorithms or employ multi-stage learning.
@@ 2.1 Knowledge Distillation
2.2 Progressive Knowledge Distillation
Building on traditional distillation methods, PKD incrementally transfers knowledge to improve the efficiency and scalability of the distillation process. In Dennis et al. (2023), the authors proposed the B-DISTIL algorithm, which decomposes a large teacher model into an ensemble of smaller student models, each capable of refining predictions as more models are evaluated (see Figure 2). This way of distillation enables a flexible trade-off between inference time and accuracy, making it particularly effective for on-device inference, where computational resources are limited.
@@ 2.2 Progressive Knowledge Distillation
Figure 2: Progressive Knowledge Distillation, the knowledge from a complex architecture teacher is split among different simple architecture students, who are trained to progressively improve a coarse prediction.
2.3 State Space Models and Selectivity
@@
2.4 Dynamic Neural Networks and Hardware-aware Parallelism
@@ 3 Theory and Methodology
3.1 Knowledge Distillation Foundation
@@
$L_{KD}=\alpha L_{hard}+(1-\alpha)L_{soft}$ (1)
@@
At a round $t$ (looking for weak learner), the algorithm maintains two probability matrices as $K^{+}_{t}\in{R}^{N\times L}$ and $K^{-}_{t}\in{R}^{N\times L}$, where $N$ is the number of data samples and $L$ is the number of labels. These matrices store the probabilities of positive and negative residual errors (differences between student and teacher predictions).
@@
$\min_{\theta}\Big(-\frac{1}{\gamma}\sum_{i,j}I_{ij}^{+}\log\Big(1+\frac{l(x_{i})_{j}}{2B}\Big)+(1-I_{ij}^{+})\log\Big(1-\frac{l(x_{i})_{j}}{2B}\Big)\Big)$ (3)
@@
3.2 Mamba Architecture Theory
@@
To make this process more efficient, as shown in (Gu et al., 2021), the SSM can be expressed as a convolution. By unrolling the recursion, the output $y_{k}$ becomes a weighted sum of the current and previous inputs, where these weights are determined by powers of the matrix $\bm{\overline{A}}$ . This leads to the construction of the convolution kernel $\bm{\overline{K}}$, which can be defined as:
$\bm{\overline{K}}$ $=(\bm{C}\bm{\overline{B}},\bm{C}\bm{\overline{A}}\bm{\overline{B}},\dots,\bm{C}\bm{\overline{A}}^{k}\bm{\overline{B}},\dots)$ (6)
$y$ $=x\ast\bm{\overline{K}}$ (7)
@@
For large sequences, the SSM processes all incoming data uniformly; this can lead to inefficiencies when much of the input data is irrelevant. Also, the representation depicted in equation 6 relies on the assumption of working in a Linear Time Invariant (LTI) System, which is not always true in complex contexts. Mamba introduces Selective-State-Space Models S6 that improve traditional SSMs by selectively processing only the most relevant parts of the input sequence. The selectivity is based on the discretization method, where the continuous $(A,B)$ are transformed to their discrete version $\overline{A_{k}},\overline{B_{k}}$ through $f_{A}(\Delta_{k},A)$ and $f_{B}(\Delta_{k},A,B)$. Rules as the zero-order hold (ZOH) defined in the following expresions can be used as $f_{A},f_{B}$ .
$\bm{\overline{A_{k}}}=\exp(\Delta_{k}\bm{A})\qquad$ (8)
$\bm{\overline{B_{k}}}=(\bm{A})^{-1}(\exp(\Delta_{k}\bm{A})-\bm{I})\cdot\bm{B}$ (9)
@@
Figure 3: Selective SSM, the flow of the data and operations, requires a hardware-aware algorithm that computes the hidden state $h$ in more efficient levels of the GPU memory hierarchy. The S-SSM is an LTVS that projects the input $x$ to construct a time dependence in $\Delta_{t},B_{t},C_{t}$. This process ensures selectivity during discretization ($A\rightarrow\overline{A},B\rightarrow\overline{B}$) and resetting the hidden state through $C_{t}$ to the final output $y_{t}$.
@@
Figure 4: Mamba Blocks inside Progressive Knowledge Distillation
3.3 Combining PKD and Mamba Architecture Methodology
@@ 4.1 Datasets
4.2 Knowledge Distillation Experiment Setup
@@ 5.1 Computational Efficiency
Table 1: MNIST overview of model configurations and some performance metrics, including the number of blocks, state dimension, total parameters, FLOPs, and accuracy for each model variant.
@@
Table 2: CIFAR-10 overview of model configurations and some performance metrics, including the number of blocks, state dimension, total parameters, FLOPs, and accuracy for each model variant.
@@
Figure 5: Accuracy vs. Number of Students
@@
Figure 6: Performance Teacher Students Comparison
5.2 Model Performance
@@
6 Conclusion and Future Work
@@
The flexibility of Mamba architecture in enabling the configuration of their hyperparameters and the introduction of more blocks improves how we construct students as weak learners. In Dennis et al. (2023), the author explained that one constraint is that the algorithm must consider pre-define a set of student classes $\mathcal{F}$}, and the way to construct these classes must be connected to the teacher architecture, expanding the architecture from a small base model. Mamba architecture, without any additional set of classes configured, allows us to propose a candidate student easily by only changing the number of blocks or slightly increasing the dimension hidden state for an SSM. The number of independent blocks in Mamba also allows for increasing complexity, going from a single-channel dataset (MNIST) to a multichannel dataset (CIFAR-10), managing the new channel sequences as new blocks of Mamba.
Additionally, future work could explore combining these techniques with Neural Architectures that are optimized to extract visual features, especially for image classification. Increasing the number of sets of classes $\mathcal{F}$} for students could relax the weak learner condition for larger and more complex datasets than those analyzed in this paper. The loss function is another characteristic that must be adapted to these newly proposed architectures. In the case of CIFAR-10 in this work, the learning condition forces the students to be more complex in their architecture, only paying attention to how this student can improve the coarse prediction. However, the concept of KD goes beyond the perception of accuracy; the KD process must consider how well the student as an individual is able to fit the knowledge from the teacher and then how well it fits its knowledge in the progressive group.
Acknowledgments